# Visual Position Prompt for MLLM based Visual Grounding

### v2

@@ IV-D Scaling Up
IV-E Transferability Studies for VPP
TABLE III: Transferability Study of VPP on LLaVA-NeXT.
@@ IV-E Transferability Studies for VPP
TABLE IV: Transferability Study of VPP-LLaVA on the ReferIt Dataset in Zero-Shot Settings
      Method       ReferIt val       ReferIt test
      LLaVA-v1.5-7B       48.95       47.42
      VPP-LLaVA-7B       57.55       56.53
Transferability studies hold two implications for our work: First, whether the proposed VPP can be transferred to models beyond LLaVA-v1.5 and similarly enhance their visual grounding capabilities; second, whether the proposed VPP-LLaVA can be transferred to other datasets in a zero-shot manner. To address these concerns, we apply the proposed VPP to LLaVA-NeXT-7B using the main experiment’s training strategy, with results in Table III. Additionally, in Table IV, we present the zero-shot visual grounding test of VPP-LLaVA-7B on the ReferIt dataset.
From Table III, we observe that the proposed VPP method can be transferred to other models, thereby enhancing their visual grounding performance.
From Table IV, it can be observed that compared to LLaVA-7B, VPP-LLaVA-7B demonstrates a significant improvement in zero-shot visual grounding performance on the ReferIt dataset. Specifically, VPP-LLaVA-7B achieves an accuracy boost of absolutely 8.6% on the ReferIt val split and 9.11% on the ReferIt test split.
IV-F Results of Region Captioning
TABLE V: Comparisons of region captioning on RefCOCOg. Models marked in gray indicate the use of larger language models. The highest performance is marked in bold, and the second-highest performance is marked with an underline.
@@ IV-F Results of Region Captioning
The language generation capability of VPP-LLaVA is also evaluated, despite not being the primary focus. Table V shows the results of region captioning on the RefCOCOg dataset. VPP-LLaVA-7B achieves a leading CIDEr score of 73.1, demonstrating comparable performance to larger models. This highlights its efficiency and ability to generate semantically rich captions. Despite its 7B parameter size and limited training data, it matches or even outperforms some 13B models in key metrics, showcasing its competitive strength. However, its METEOR score of 12.1 indicates potential room for improvement in caption accuracy. For other MLLMs, as seen in many published works such as Groma, LISA, and Griffon, general language capabilities are often compromised due to the size and type of the SFT datasets. This remains a significant challenge in the MLLM field, which we leave for future work.
@@ IV-G Ablation Study
TABLE VI: The ablation study of different modules in our proposed method.
@@
To conserve time and computational resources, we freeze the visual encoder (except in Section IV-G6) and randomly select 150K samples from VPP-SFT to form our ablation training set. All subsequent ablation results are based on this subset and are specific to the VPP-LLaVA-7B model.
@@ IV-G1 Components
Table VI shows VPP-LLava’s performance with different components on grounding benchmarks. Without the proposed VPP, the model degrades to LLaVA-v1.5 (the baseline, which is also trained on the same 150K subset with the same training strategy as described earlier for fairness). Adding the global VPP improves baseline performance across datasets, especially on RefCOCOg, with a 1.5% gain, highlighting its effectiveness in providing a reliable spatial reference. The local VPP alone shows mixed results, sometimes comparable to or slightly below baseline, likely due to limited data hindering effective DETR feature alignment. However, combining global and local VPPs yields a synergistic effect, enhancing grounding accuracy by around 2% on RefCOCO testB, as the global reference supports overall spatial orientation while the local VPP contributes region-specific semantic cues, jointly boosting performance.
@@ IV-G2 Impact of Trainable Global VPP
TABLE VII: The ablation study of the impact about the global VPP being trainable or non-trainable.
@@
In our approach, the default setting allows the global VPP to be trainable. It is essential to examine the impact of training versus not training the global VPP. To conserve time and computational resources, we maintain the same setup as in our previous ablation study, using a 150K subset sampled from VPP-SFT while keeping CLIP frozen. The results are presented in Table VII. We can observe that making the global VPP learnable is beneficial. This is because the pixel values of the axis-like images we use for initialization may not be suitable for every image. By making it learnable, the model can dynamically adapt to different inputs, thereby achieving better performance.
@@ IV-G3 Variants of Global VPP Initialization
TABLE VIII: The ablation study on variants of global VPP Initialization.
@@
The Global VPP is initialized by default with an axes image that features coordinate axes along its edges, using a unit scale of 0.1. To evaluate how different initial axes images affect the grounding performance of the model, we conduct experiments under two distinct conditions: one with a unit scale of 0.05 and axes along the edges (Interval-0.05 in Table VIII), and another with a unit scale of 0.1 and cross axes centered at the image (Cross-axis-0.1 in Table VIII). It should be noted that when the cross axes are equipped, the binary mask needs to be removed. The results show that the default settings that the global VPP around all edges with the 0.1 unit scale, as used in our paper, performs better. Furthermore, the interval should be based on the internal coordinate system of the MLLM base model due to differences among MLLMs. We will explore more initialization forms in the future.
IV-G4 Type of Local VPP Generator
TABLE IX: The ablation study of different types of local VPP generator. A.DETR and V.DETR are the abbreviation for anchor-DETR and vanilla-DETR respectively.
@@ IV-G4 Type of Local VPP Generator
We compare the performance of using different types of DETR to obtain the local VPP, including anchor-DETR (A.DETR) [16] and vanilla-DETR (V.DETR) [15]. As shown in Table IX, A.DETR is generally inferior to V.DETR. This may be due to the excessive number of object queries (900 for A.DETR and 100 for V.DETR) interfering with the LLM’s judgment. Furthermore, it remains to be investigated in future work whether a more powerful DETR variant could improve visual grounding performance.
IV-G5 Fusion Strategy
TABLE X: The ablation study of the fusion strategy. C.A.1, C.A.2 and Cat. are the abbreviation for cross-attention-1, cross-attention-2, and concatenate operation respectively.
@@ IV-G5 Fusion Strategy
Table X presents our ablation study on fusion strategies for combining $F_{gp}^{\prime}$ (CLIP visual features with global VPP) and $F_{lp}^{\prime}$ (features from the local VPP generator) as outlined in Eq. 8. Here, C.A.1 stands for cross-attention-1, which uses $F_{lp}^{\prime}$ as the query and $F_{gp}^{\prime}$ as the key in cross-attention, while C.A.2 reverses this, using $F_{gp}^{\prime}$ as the query and $F_{lp}^{\prime}$ as the key. Cat. denotes direct concatenation. The cross-attention modules, trained with a learning rate of 2e-4, underperform relative to simple concatenation. This may be due to their training from scratch, which limits alignment with the LLM’s pre-trained representations. Consequently, we adopt direct concatenation for our final model.
IV-G6 Freeze or Unfreeze Visual Encoder
TABLE XI: The ablation study examines whether to freeze or unfreeze the visual encoder.
@@ IV-G6 Freeze or Unfreeze Visual Encoder
In Table XI, we report the results of freezing versus unfreezing the MLLM’s visual encoder in our models. We observe that unfreezing the visual encoder significantly enhances performance on downstream visual grounding task. Since the introduced global VPP does not naturally exist in the MLLM’s pretraining dataset, unfreezing the visual encoder allows it to better learn and adapt to this new form of input. This observation is also supported by recent research [21, 68, 4].
IV-G7 Hyperparameters
@@ V-A Qualitative Dialogue
To fully understand the position-enhanced VPP-LLaVA for grounding tasks, we present our qualitative dialogue in Fig. 6. The predicted bounding box is marked in red, and the ground truth is marked in green. For brevity, some instruction prompts are omitted, and the full form of the conversation can be found in Table I. Specifically, cases (a), (b), and (c) are three visual grounding examples from RefCOCO, RefCOCO+, and RefCOCOg, respectively. We can observe that our VPP-LLaVA demonstrates strong visual grounding capacity on these widely used benchmarks through the proposed global and local VPP.
@@ V-B Qualitative Comparisons
V-C Failure Cases Study

### v4

@@ IV-D Scaling Up
IV-E Zero-Shot Generalization on Unseen Datasets
TABLE III: Zero-Shot Generalization Comparisons on ReferIt Dataset.
      Method       ReferIt val       ReferIt test
      LLaVA-v1.5-7B       48.95       47.42
      VPP-LLaVA-7B       57.55       56.53
TABLE IV: Zero-shot generalization results on the newly proposed challenging GSEval-BBox benchmark. Our method achieves a significant performance lead, demonstrating strong generalization to unseen grounding scenarios.
Method GSEval-BBox
Stuff Part Multi Single All
InternVL2.5-78B [80] 85.3 16.8 63.2 55.7 62.2
InternVL2.5-8B [80] 91.3 7.3 65.7 47.3 58.2
Qwen2.5-VL-72B [34] 88.4 31.2 42.8 64.6 62.5
Qwen2.5-VL-7B [34] 93.0 17.6 75.9 59.1 66.7
DeepSeek-VL2 [81] 86.5 12.7 64.8 51.2 60.8
Mistral-Small-3.1-24B [82] 16.0 3.5 20.3 10.7 13.2
Ferret-7B [7] 82.1 17.6 56.0 43.2 53.3
Ferret-13B [7] 80.6 21.3 58.0 46.6 55.1
VPP-LLaVA-7B 52.1 67.9 82.7 68.7 67.0
VPP-LLaVA-13B 54.5 69.9 83.8 69.5 68.4
We evaluate the zero-shot generalization capability of VPP-LLaVA on two datasets not seen during training: ReferIt and the newly proposed GSEval-BBox. Since ReferIt has been included in the training set of many existing MLLMs, we report only the performance of the baseline LLaVA-v1.5-7B for fair comparison. As shown in Table III, VPP-LLaVA-7B outperforms the baseline by a large margin, with an absolute accuracy gain of 8.6% on the validation split and 9.1% on the test split.
To further assess the robustness and generalization of our approach in complex scenarios, we conduct a comprehensive zero-shot evaluation on GSEval-BBox [33], a challenging benchmark featuring diverse expression types including stuff, part-object, multi-object, and single-object. As shown in Table IV, VPP-LLaVA-13B achieves 69.9% accuracy on the part split, significantly outperforming the next best model (Qwen2.5-VL-72B) by a wide margin. On multi-object expressions, VPP-LLaVA-7B attains 82.7%, clearly ahead of Qwen2.5-VL-7B. For single-object cases, our model also achieves the highest score of 69.5%, setting a new state-of-the-art.
These results highlight the strong generalization ability of VPP-LLaVA in zero-shot visual grounding. We attribute this performance to the synergy between our proposed VPP design and the high-quality supervision provided by the VPP-SFT dataset, which enables the model to generalize well across diverse and challenging scenarios.
IV-F Transferability Studies for VPP
TABLE V: Transferability Study of VPP on LLaVA-NeXT.
@@ IV-F Transferability Studies for VPP
TABLE VI: Comparisons of region captioning on RefCOCOg. Models marked in gray indicate the use of larger language models. The highest performance is marked in bold, and the second-highest performance is marked with an underline.
@@
Transferability studies in our framework focus on two key aspects: (1) whether the proposed VPP mechanism can be effectively transferred to MLLMs beyond LLaVA-v1.5; and (2) whether VPP-LLaVA preserves strong language generation capabilities.
For question (1), we apply VPP to LLaVA-NeXT-7B with the same training strategy, and the results in Table V show consistent improvements in visual grounding performance. This confirms that VPP effectively enhances grounding ability across different model architectures.
For question (2), we evaluate the language generation capability of VPP-LLaVA on the coordinate-aware language generation task, i.e., region captioning, using the RefCOCOg dataset. As shown in Table VI, VPP-LLaVA-7B achieves a CIDEr score of 73.1, comparable to or better than some larger models. Despite its smaller size and relatively limited training data, VPP-LLaVA maintains strong localization and captioning performance.
@@ IV-G Ablation Study
TABLE VII: The ablation study of different modules in our proposed method.
@@
To conserve time and computational resources, we freeze the visual encoder (except in Section IV-G7) and randomly select 150K samples from VPP-SFT to form our ablation training set. All subsequent ablation results are based on this subset and are specific to the VPP-LLaVA-7B model.
@@ IV-G1 Components
Table VII shows VPP-LLava’s performance with different components on grounding benchmarks. Without the proposed VPP, the model degrades to LLaVA-v1.5 (the baseline, which is also trained on the same 150K subset with the same training strategy as described earlier for fairness). Adding the global VPP improves baseline performance across datasets, especially on RefCOCOg, with a 1.5% gain, highlighting its effectiveness in providing a reliable spatial reference. The local VPP alone shows mixed results, sometimes comparable to or slightly below baseline, likely due to limited data hindering effective DETR feature alignment. However, combining global and local VPPs yields a synergistic effect, enhancing grounding accuracy by around 2% on RefCOCO testB, as the global reference supports overall spatial orientation while the local VPP contributes region-specific semantic cues, jointly boosting performance.
@@ IV-G2 Impact of Trainable Global VPP
TABLE VIII: The ablation study of the impact about the global VPP being trainable or non-trainable.
@@
In our approach, the default setting allows the global VPP to be trainable. It is essential to examine the impact of training versus not training the global VPP. To conserve time and computational resources, we maintain the same setup as in our previous ablation study, using a 150K subset sampled from VPP-SFT while keeping CLIP frozen. The results are presented in Table VIII. We can observe that making the global VPP learnable is beneficial. This is because the pixel values of the axis-like image used for initialization may not be optimal for all images. By making the global VPP learnable, the model adjusts its parameters during training in a data-driven manner to better encode position-related cues. Although the global VPP remains fixed during inference, the learned representation provides generally useful spatial cues across inputs.
@@ IV-G3 Variants of Global VPP Initialization
TABLE IX: The ablation study on variants of global VPP Initialization.
@@
To evaluate how different initializations affect the model’s grounding performance, we conduct experiments under four distinct conditions, as shown in Fig. 3: (1) using a unit scale of 0.1 with axes along the edges (Default in Table IX); (2) using a unit scale of 0.05 with axes along the edges (Internal-0.05 in Table IX); (3) using a unit scale of 0.1 with cross-shaped axes centered in the image (Cross-axis-0.1 in Table IX); and (4) using a unit scale of 0.1 with externally padded axes along the edges (External-0.1 in Table IX). Note that for conditions (3)) and (4), the binary mask must be removed. We can draw several observations as follows. First, the results show that the default setting achieves better performance. Second, the External-0.1 strategy introduces a scale mismatch, as the actual image content occupies only 276×276 pixels within the 336×336 input. Since the pre-processing in our baseline model and the training data is based on the full 336×336 coordinate space, this mismatch leads to degraded performance. Lastly, the interval scale should be aligned with the internal coordinate system of the base MLLM, as different MLLMs may have varying internal representations.
IV-G4 Font Size of Global VPP Initialization
TABLE X: The ablation study on font size of the axis-like image for global VPP initialization.
Font Size RefCOCO RefCOCO+ RefCOCOg
val testA testB val testA testB val test
5 85.05 89.94 79.69 77.47 84.03 67.70 79.92 79.73
15 84.90 89.39 79.68 77.46 83.74 67.05 79.11 79.38
Default (10) 85.29 89.96 80.06 77.68 84.75 67.62 79.80 80.36
To further analyze the effect of initialization on model performance, we investigate the impact of font size used for coordinate digits in the axis-like image that initializes the global VPP. As shown in Table X, a moderate font size achieves the best trade-off. Both excessively small and large fonts result in slight performance degradation. Based on empirical observations, we adopt a font size of 10, which prevents digit overlap and ensures clear visual boundaries.
IV-G5 Type of Local VPP Generator
TABLE XI: Ablation study of different local VPP generators. A.DETR and V.DETR refer to Anchor-DETR and Vanilla-DETR. denotes our adapted version (100 queries for A.DETR and 116 queries for YOLO-World).
@@ IV-G5 Type of Local VPP Generator
We further investigate the effect of using different detectors to generate local VPP features. As shown in Table XI, we compare Anchor-DETR (A.DETR) [16], Vanilla-DETR (V.DETR) [15], and the open-vocabulary detector YOLO-World [83].
V.DETR consistently yields the strongest performance across all splits. Interestingly, reducing the number of object queries in A.DETR improves its results, suggesting that an excessive number of proposals may interfere with the LLM’s spatial reasoning. Meanwhile, the YOLO-World variant—where we adapt its multi-scale features via downsampling and project them into the LLM space—also performs competitively, demonstrating the flexibility of our method to work with both closed-set and open-vocabulary detectors.
Moreover, as shown in our previous zero-shot evaluations on ReferIt and GSEval-BBox, our default V.DETR setup already exhibits strong zero-shot/open-vocabulary capabilities when integrated into the VPP framework. For these reasons, we opt to use V.DETR in our main experiments.
IV-G6 Fusion Strategy
TABLE XII: The ablation study of the fusion strategy. C.A.1, C.A.2 and Cat. are the abbreviation for cross-attention-1, cross-attention-2, and concatenate operation respectively.
@@ IV-G6 Fusion Strategy
Table XII presents our ablation study on fusion strategies for combining $F_{gp}^{\prime}$ (CLIP visual features with global VPP) and $F_{lp}^{\prime}$ (features from the local VPP generator) as outlined in Eq. 8. Here, C.A.1 stands for cross-attention-1, which uses $F_{lp}^{\prime}$ as the query and $F_{gp}^{\prime}$ as the key in cross-attention, while C.A.2 reverses this, using $F_{gp}^{\prime}$ as the query and $F_{lp}^{\prime}$ as the key. Cat. denotes direct concatenation. The cross-attention modules, trained with a learning rate of 2e-4, underperform relative to simple concatenation. This may be due to their training from scratch, which limits alignment with the LLM’s pre-trained representations. Consequently, we adopt direct concatenation for our final model.
IV-G7 Freeze or Unfreeze Visual Encoder
TABLE XIII: The ablation study examines whether to freeze or unfreeze the visual encoder.
@@ IV-G7 Freeze or Unfreeze Visual Encoder
In Table XIII, we report the results of freezing versus unfreezing the MLLM’s visual encoder in our models. We observe that unfreezing the visual encoder significantly enhances performance on downstream visual grounding task. Since the introduced global VPP does not naturally exist in the MLLM’s pretraining dataset, unfreezing the visual encoder allows it to better learn and adapt to this new form of input. This observation is also supported by recent research [22, 76, 4].
IV-G8 Impact of VPP-Related Text Instructions
TABLE XIV: The Ablation Study on Levels of VPP-related Text Instructions.
System Sample RefCOCO RefCOCO+ RefCOCOg
val testA testB val testA testB val test
85.17 89.37 79.67 77.02 84.58 67.25 78.92 79.59
85.19 90.03 80.00 77.98 84.54 67.72 79.53 79.47
85.29 89.96 80.06 77.68 84.75 67.62 79.80 80.36
By default, each training sample in VPP-SFT includes a VPP-related text instruction at the sample level, i.e., Each image is accompanied by axes. If the question pertains to the bounding box coordinates, refer to the axes for the response. To assess the impact of this instruction on model performance, we compare three variants, as reported in Table XIV. The first variant removes the instruction entirely during training and testing. The second places the instruction at the system level (i.e., system prompt) but omits it at the sample level. The third follows our default setting, where the instruction is included at the sample level.
Experimental results indicate that removing VPP-related text instructions at both the system and sample levels leads to a consistent drop in performance, particularly on semantically complex benchmarks such as RefCOCOg. This suggests that explicit textual guidance is essential for effectively leveraging the global VPP. While the performance difference between system-level and sample-level instruction is relatively small, system-level prompts introduce global constraints that may hinder the model’s generalization to other tasks or instruction formats. To retain flexibility and task-specific adaptability, we adopt sample-level instruction as the default.
IV-G9 Hyperparameters
@@ V Qualitative Results
Figure 10: Qualitative visual grounding results of our VPP-LLaVA compared with the state-of-the-art method Qwen2.5-VL-7B on GSEval-BBox dataset with zero-shot settings. For brevity, some text instructions are omitted. Green: ground truth; Red: ours; Purple: Qwen2.5-VL-7B.
@@ V-A Qualitative Dialogue
To fully understand the position-enhanced VPP-LLaVA for grounding tasks, we present our qualitative dialogue in Fig. 8. The predicted bounding box is marked in red, and the ground truth is marked in green. For brevity, some instruction prompts are omitted, and the full form of the conversation can be found in Fig. 5. Specifically, cases (a), (b), and (c) are three visual grounding examples from RefCOCO, RefCOCO+, and RefCOCOg, respectively. We can observe that our VPP-LLaVA demonstrates strong visual grounding capacity on these widely used benchmarks through the proposed global and local VPP.
@@ V-B Qualitative Comparisons on RefCOCO/+/g
V-C Qualitative Comparisons on Zero-shot GSEval-BBox
To further evaluate the robustness of our VPP-LLaVA model in zero-shot scenarios, we present qualitative results on the GSEval-BBox dataset. As a comparison method, we include Qwen2.5-VL-7B, which ranks second in Table IV and represents one of the most advanced open-source MLLMs. The bounding boxes predicted by Qwen2.5-VL-7B are shown in purple. We illustrate several challenging referent localization cases across different levels of granularity.
From the results, we observe several key findings. In conjunction with Table IV, although our model shows a performance gap on stuff categories—likely due to limited data coverage and possible dataset bias—we still achieve competitive overall results, even under complex textual prompts. Notably, our model demonstrates remarkable performance on part-object cases. This is particularly evident in cases (b) and (h), where small object recognition poses a long-standing challenge in visual grounding, especially when the target is a part of a larger object and the prompt is linguistically complex. In these cases, our model produces highly accurate localizations. In contrast, even the state-of-the-art Qwen2.5-VL-7B, despite being trained on significantly more data, fails to precisely identify the part-object and instead predicts bounding boxes covering the entire main object—for example, the full motorcycle in case (b) and the whole apple in case (h).
Moreover, our model consistently handles both multi-object and single-object scenarios well on this dataset. While Qwen2.5-VL-7B also performs reasonably in some of these cases, a performance gap remains in favor of VPP-LLaVA. For instance, in case (i), our model accurately localizes two pizzas as instructed by the prompt, whereas Qwen2.5-VL-7B identifies only one. In case (e), Qwen2.5-VL-7B even fails to recognize the correct referent, mistakenly localizing an unrelated object, while our model identifies the intended target precisely.
These visualizations highlight the effectiveness of our VPP-LLaVA model, which integrates both global and local positional references. This combination is particularly beneficial in complex scenes where spatial understanding is crucial. The strong performance also stems from our high-quality VPP-SFT data—despite a relatively small training scale of 0.6M samples, our model achieves impressive zero-shot results, especially on challenging expressions and referent localization. These observations further demonstrate the robustness and efficiency of our approach.
V-D Failure Cases Study
@@ VI Conclusion and Limitations
The main limitations of this work are as follows. First, although our method achieves strong gains with limited training data, it still encounters challenges in handling complex relational reasoning and queries involving ambiguity or negation, which require deeper contextual understanding. Second, while the model shows strong performance and generalization in coordinate-related tasks, its capabilities on broader multimodal tasks remain underexplored and present opportunities for further enhancement. Future work may focus on incorporating more diverse and large-scale supervision or reinforcement learning methods to improve general MLLM abilities.
VII Acknowledgments
This work was supported by National Natural Science Foundation of China (Grant No. 62425603), Basic Research Program of Jiangsu Province (Grant No. BK20240011), and Shanghai Artificial Intelligence Laboratory.